A survey asked, "How many tattoos do you currently have on your body?" Of the 1207 males surveyed, 187 responded that they had at least one tattoo. Of the 1098 females surveyed, 133 responded that they had at least one tattoo.

Construct a 90 % confidence interval to judge whether the proportion of males that have at least one tattoo differs significantly from the proportion of females that have at least one tattoo. Interpret the interval.

Let p1 represent the proportion of males with tattoos and p2 represent the proportion of females with tattoos. Find the 90 % confidence interval for p1 - p2.

The lower bound is:

The upper bound is:
Answer #1

p1cap = 187/1207 = 0.1549
p2cap = 133/1098 = 0.1211

pcap = (187 + 133)/(1207 + 1098) = 0.1388

SE = sqrt(0.1388*(1-0.1388)*(1/1207 + 1/1098))
SE = 0.0144

for 90% CI, z-value = 1.645

lower bound = 0.1549 - 1.645*0.0144 = 0.1312
upper bound = 0.1549 + 1.645*0.0144 = 0.1786